# Style-Friendly SNR Sampler for Style-Driven Generation

### v2

@@ Abstract
Recent large-scale diffusion models generate high-quality images but struggle to learn new, personalized artistic styles, which limits the creation of unique style templates. Fine-tuning with reference images is the most promising approach, but it often blindly utilizes objectives and noise level distributions used for pre-training, leading to suboptimal style alignment. We propose the Style-friendly SNR sampler, which aggressively shifts the signal-to-noise ratio (SNR) distribution toward higher noise levels during fine-tuning to focus on noise levels where stylistic features emerge. This enables models to better capture unique styles and generate images with higher style alignment. Our method allows diffusion models to learn and share new “style templates”, enhancing personalized content creation. We demonstrate the ability to generate styles such as personal watercolor paintings, minimal flat cartoons, 3D renderings, multi-panel images, and memes with text, thereby broadening the scope of style-driven generation.
Figure 1: Our method learns style templates from reference images, capturing elements including color schemes, layouts, illumination, and brushstrokes. For each image, we fine-tune diffusion models using the reference image in the red insert box to generate the output. We show a ‘fluffy baby sloth with a knitted hat trying to figure out a laptop’ in different meme templates: a meme with the words ‘you just activated my trap card’ (top left), a multi-panel comic layout (bottom left), and a two-panel meme (top middle). On the right side, we display typographies in the styles of a wooden sculpture and a minimal line drawing. We also present a ‘singing kangaroo drinking beer’ in various artistic styles—flat illustration, crayon drawing, watercolor, and line drawing.
@@ 1 Introduction
Recently, large-scale text-to-image diffusion models [42, 8, 38, 37, 47, 29, 1] have achieved remarkable progress in visual content creation. In particular, open-weights like Stable Diffusion series [42, 8] and FLUX [29] have been among the most notable for their photorealistic image quality and language understanding capabilities. Behind this strong performance lies the advancement of the diffusion framework that encompasses the principles of score-based models [55] and flow matching [32, 33], diffusion formulations [55, 32, 33, 17], loss weighting [4, 24], noise level scheduling [18, 22], and architectural improvements [36, 8, 23]. These advancements have predominantly focused on generating high-quality images with respect to object-centric benchmarks [21, 11] and metrics [15, 28].
@@
Motivated by the success of text-to-image models, there is a growing need for style-driven generation, where the generated samples capture styles desired by individual users or artists. Here, “style” encompasses various elements such as color schemes, layouts, illumination, and brushwork [10, 51, 31, 7]—all contributing to the unique nuances of an image. However, relying solely on text prompts has its limitations in capturing these complex and personal artistic styles, especially those not present in the pre-training data.
@@
To enable style-driven generation, researchers have turned to fine-tuning [45, 51] text-to-image models [47, 3] using reference images of the desired styles, often employing lightweight adapters such as LoRA [19], as this remains the most effective method for accurately capturing and reproducing new artistic styles. However, they blindly apply the same objective functions and noise level distributions used for pre-training—originally optimized for object-centric benchmarks [11, 21]—without considering the unique characteristics inherent to style images, which are distinct from object-centric images. Consequently, even with numerous style-driven generation studies [51, 30, 61, 44, 14], we observe prevalent failure cases when fine-tuning diffusion models on style references. Addressing these failures sometimes requires two-stage training with heavy human feedback [51].
In this paper, we address these limitations by introducing the Style-friendly SNR sampler, a method that significantly enhances the model’s ability to capture new artistic styles during fine-tuning. Our approach is motivated by two key observations 1) diffusion models struggle to learn new styles, and 2) styles emerge at higher noise levels. Building upon these observations, we propose adjusting the noise level sampling in the diffusion model’s objective function during fine-tuning. By directly sampling signal-to-noise ratio (SNR) using our Style-friendly SNR sampler, we bias the distribution towards higher noise levels where style features emerge, such as color schemes, layouts, and illumination—key elements for style representation. Unlike previous methods [8, 30], our approach intensifies the focus on capturing these stylistic aspects.
@@
Our Style-friendly SNR sampler enables state-of-the-art models such as FLUX-dev [29] and Stable Diffusion 3.5 [8, 57] to effectively learn style templates with unprecedented accuracy. Our method leads to a more faithful style-driven generation, capturing the uniqueness of reference style images. Moreover, we unveil the key components that explain why previous works excel at learning object-centric concepts but struggle with styles, providing deeper insights into the diffusion process for style-driven generation. Ultimately, our approach facilitates the creation of style templates from reference images, which can be easily shared and utilized to practitioners for content creation, expanding the capabilities of text-to-image diffusion models.
@@ 2 Training Diffusion Models
Figure 2: Fine-tuning capability. While FLUX succeeds in learning objects (a), it struggles to capture styles (b). We enable FLUX to learn styles (c). References are shown in the red insert box.
Figure 3: Prompt switching during generation. $\lambda_{t}$ indicates log-SNR. The style prompts are ‘minimalist flat round logo’, ‘sticker’, ‘detailed pen and ink drawing’, and ‘abstract rainbow colored flowing smoke wave’. Styles emerge in the initial 10% of denoising steps; therefore, (c) and (d) fail to capture target styles. Here, we use FLUX, with the guidance scale 7 across the whole denoising process.
2.1 Diffusion Process
@@
where $\alpha_{t}$ and $\sigma_{t}$ are predefined noise schedules, and $\epsilon\sim\mathcal{N}(0,I)$ represents standard Gaussian noise. Recent state-of-the-art diffusion models, such as Stable Diffusion 3 (SD3) [8] and FLUX [29], utilize the noise schedule from rectified flow [32, 33] with $\alpha_{t}=1-t$ and $\sigma_{t}=t$. This choice is effective due to straight diffusion trajectories.
While the diffusion process is commonly parameterized by the timestep $t$, Kingma et al. [25, 24] characterize the noise level using the log signal-to-noise ratio (log-SNR):
@@
In the case of rectified flow, $\lambda_{t}=2\log\left(\frac{1-t}{t}\right)$. In our work, we demonstrate that adjusting the timestep distribution based on $\lambda_{t}$ facilitates learning styles more effectively than using $t$.
2.2 Unified Loss Function
Kingma et al. [25, 24] showed that the objectives of various diffusion models [17, 42, 32, 33, 22, 54, 55] share the same form. The general loss function is expressed as:
$\mathcal{L}_{DM}(x_{0})=\mathbb{E}_{\lambda\sim p(\lambda)}\bigg{[}||f_{\theta}(x_{\lambda},\lambda)-f(x_{\lambda},\lambda)||^{2}\bigg{]},$ (3)
where $f(x_{\lambda},\lambda)$ is the target function. In the case of rectified flow, the model $f_{\theta}$ predicts the velocity $f(x_{\lambda},\lambda)=\epsilon-x_{0}$.
Considering the role of $p(\lambda)$, it serves as a weighting factor in the loss function that determines the importance assigned to different noise levels during training:
[…]
$=\mathbb{E}_{\lambda\sim U(\lambda_{\text{min}},\lambda_{\text{max}})}\bigg{[}p(\lambda)||f_{\theta}(x_{\lambda},\lambda)-f(x_{\lambda},\lambda)||^{2}\bigg{]}.$
From this perspective, we can observe the inductive bias imposed on the diffusion model through $p(\lambda)$. Prior works [4, 6] discuss that the model learns low-frequency information at large noise levels (small $\lambda$) and learns high-frequency details at small noise levels (large $\lambda$). Consequently, diffusion models generate images in a coarse-to-fine manner [41, 6].
2.3 SNR Samplers
Previous diffusion models have predominantly been designed based on the timestep $t$ rather than the log-SNR $\lambda_{t}$, with $t$ typically sampled from a uniform distribution [17, 42, 36] during training. Thus, the importance of each noise level in diffusion loss is determined solely by the function $\lambda_{t}$. However, EDM [22] argues that this approach entangles the design space of diffusion models and proposes sampling noise variance $\sigma_{t}$ from a log-normal distribution.
SD3 [8] extends this approach to the training of text-to-image models by sampling logit function of timestep $\log\left(\frac{t}{1-t}\right)$ from a normal distribution. As mentioned earlier, SD3 follows the rectified flow formulation and has the log-SNR as $\lambda_{t}=2\log\left(\frac{1-t}{t}\right)$. In addition, they propose shifting timestep $t$ to $t_{\text{new}}$ by $k$ for high resolution training:
@@ 2.3 SNR Samplers
$\lambda_{t_{\text{new}}}=2\log\left(\frac{1-t_{\text{new}}}{t_{\text{new}}}\right)=\lambda_{t}-2\log k.$ (6)
@@ 3.1 Observations
We begin by examining the fine-tuning capability of a recent state-of-the-art model. Many current text-to-image models concentrate on object-centric image quality and demonstrate their performance. Among the latest diffusion models, FLUX [29], the representative text-to-image model, excels in object-driven image generation through fine-tuning, enabling high-quality generation of fine-tuned objects across various scenarios. In Fig. 2a, the dog-themed backpack integrates well into the nighttime surfing scene, matching the appropriate evening lighting. The dog below is dressed in new clothes, and the typography is also well-generated. However, FLUX struggles to capture styles in terms of color schemes, illumination, and brushwork when fine-tuning with the SNR sampler used for pre-training [8]. In Fig. 2b, the “glowing” example only causes the sloth’s fur to glow but does not reflect the dark background and blue lighting of the reference. Similarly, for the Van Gogh oil painting reference, while it captures the blue color tone, it fails to reproduce Van Gogh’s distinctive brushstrokes.
@@
To investigate the failure of fine-tuning styles, we perform further analysis related to styles. We investigate the noise level at which the style features emerge when generating images with text-to-image diffusion models. We incorporate style descriptions, such as ‘in flat cartoon illustration’ only at specific intervals during the denoising process.
As illustrated in Fig. 3, we implement the text-conditioned generation process using text $y_{\text{w/o style}}$ to exclude style descriptions during the early denoising steps, and then switch to text $y_{\text{w/ style}}$ to include style descriptions in the remaining steps. When style descriptions are used throughout all denoising steps, the generated images align well with desired styles (Fig. 3a). Interestingly, omitting style descriptions during just the initial 10% of the denoising steps results in images where styles are not adequately reflected (Fig. 3c). These images closely resemble those generated without any style descriptions at all—yielding photorealistic samples (Fig. 3d). The case that uses $y_{\text{w/o style}}$ in the later denoising steps is provided in Fig. S8.
This observation indicates that the stylistic features are determined during the early stage of the denoising process, corresponding to an interval where the log-SNR $\lambda_{t}$ is small (i.e., higher noise levels). The subsequent denoising steps primarily render content and fine details that are independent of styles. Thus, missing the style prompt in the initial steps significantly hinders the model’s ability to generate images in desired styles, emphasizing the importance of early-stage conditioning for style representation.
@@ 3.2 Style-Friendly SNR Sampler
Figure 4: Probability distribution of Log-SNR. Colored region indicates style-emerging noise levels discussed in Sec. 3.1.
Figure 5: Effect of varying $\mu$. Diffusion models start to capture the reference glowing style when $\mu$ is lower than $-4$. The target prompt is ‘A Christmas tree in glowing style’.
@@
Figure 6: SNR sampler analysis. DINO similarities of varying SNR sampler parameters with FLUX and SD3.5-8B. Dotted lines in (c) indicate results of SD3 sampler [8]. Unless specified, we use $\mu=-6$, $\sigma=2$, and rank 32. CLIP scores are shown in Fig. S7.
Our observations indicate the key motivation for style learning. While styles emerge in the early steps of the denoising process, the current fine-tuning process utilizes an SNR sampler from pre-training which prioritizes object-centric generation [21, 11] as shown in the green line of Fig. 4. This SNR sampler places greater emphasis on intermediate steps to capture fine details of objects better, yet it does not sufficiently focus on the noise levels where styles emerge. As a result, despite excelling in object-driven generation, the current fine-tuning struggles to fully capture and represent target styles in style-driven generation.
$\lambda_{t}\sim\mathcal{N}(\mu_{\text{low}},\sigma^{2}),$ (7)
with a lowered mean $\mu_{\text{low}}$, thereby focusing the training on the critical noise levels essential for style learning. We set the standard deviation $\sigma=2$ to maintain a balance between concentrating on the critical noise levels and providing sufficient variation in noise levels for effective learning.
While timestep shifting in Eq. 5 weakly biases the noise level distribution, a Style-friendly SNR sampler allows for a more aggressive emphasis on desired noise levels. Setting the mean to $\mu=-6$ concentrates timesteps between 0.8 and 1.0, where style features emerge (Sec. 3.1). This enables us to support the diffusion model in fine-tuning with a strong focus on styles across various style templates.
3.3 Trainable Parameters of MM-DiT
We fine-tune both FLUX-dev [29] and SD3.5 [8, 57] by training LoRA [19] adapters on specific layers to capture new styles. Currently, no studies have fine-tuned the Multi-Modal Diffusion Transformer (MM-DiT) [8], the core architecture for both FLUX-dev and SD3.5. MM-DiT comprises dual-stream transformer blocks with separate parameters for text tokens and image tokens, which interact through joint attention mechanisms. To effectively learn the stylistic features encompassing both visual and linguistic characteristics, we train LoRA on the attention layers of both modalities. Additionally, FLUX includes single-stream blocks that handle both modalities simultaneously with attention mechanisms and projection layers that skip this attention, to which we also apply LoRA. This targeted fine-tuning achieves high style-alignment without training the entire network, providing a parameter-efficient method for fine-tuning MM-DiT.
@@ 4 Experiments
We evaluate our method by fine-tuning FLUX-dev [29] and SD3.5-8B [8, 57] on 18 reference styles from the StyleDrop [51]. For each reference style, we generate two images for each of the 23 evaluation prompts collected from [51], resulting in a total of 828 images evaluated per experiment.
For quantitative evaluation, we assess style alignment using DINO [2] ViT-S/16 and CLIP [39] ViT-B/32 image similarity (CLIP-I), and we measure alignment to the target prompts using CLIP text-image similarity (CLIP-T).
All models are fine-tuned using the Adam optimizer [26] for 300 steps at a learning rate of $10^{-4}$. We perform LoRA [19] fine-tuning on the frozen pre-trained models, using a rank of 32 in all experiments except for the rank ablation studies. During inference, we use 28 denoising steps and set the guidance [16, 34] scale to 7.0.

### v3

@@ Abstract
Recent text-to-image diffusion models generate high-quality images but struggle to learn new, personalized styles, which limits the creation of unique style templates. In style-driven generation, users typically supply reference images exemplifying the desired style, together with text prompts that specify desired stylistic attributes. Previous approaches popularly rely on fine-tuning, yet it often blindly utilizes objectives and noise level distributions from pre-training without adaptation. We discover that stylistic features predominantly emerge at higher noise levels, leading current fine-tuning methods to exhibit suboptimal style alignment. We propose the Style-friendly SNR sampler, which aggressively shifts the signal-to-noise ratio (SNR) distribution toward higher noise levels during fine-tuning to focus on noise levels where stylistic features emerge. This enhances models’ ability to capture novel styles indicated by reference images and text prompts. We demonstrate improved generation of novel styles that cannot be adequately described solely with a text prompt, enabling the creation of new style templates for personalized content creation.
Figure 1: Fine-tuning text-to-image diffusion models on the style-friendly noise levels enables learning novel styles from reference images and text prompts. We present ‘A kangaroo holding a beer, wearing ski goggles and passionately singing silly songs’ in various styles including watercolor painting, flat illustration, and 3d rendering styles. References are shown in the red insert box.
@@ 1 Introduction
Recently, large-scale text-to-image diffusion models [44, 9, 41, 40, 49, 31, 2] have achieved remarkable progress in visual content creation. In particular, open-weights such as Stable Diffusion series [44, 9] and FLUX [31] have been among the most notable for their photorealistic image quality and language understanding capabilities. Behind this strong performance lies the advancement of the diffusion framework that encompasses the principles of score-based models [57] and flow matching [35, 36], diffusion formulations [57, 35, 36, 18], loss weighting [5, 26], noise level scheduling [19, 24], and architectural improvements [39, 9, 25]. These advancements have predominantly focused on generating high-quality images with respect to object-centric benchmarks [22, 12] and metrics [16, 30].
@@
Motivated by the success of text-to-image models, there is a growing need for style-driven generation [53, 15, 46, 62, 32], where the generated samples capture styles desired by individual users or artists. Here, “style” encompasses various elements such as color schemes, layouts, illumination, and brushwork [11, 53, 33, 8], all contributing to the unique nuances of an image, and is typically specified or clarified using text prompts. However, relying solely on prompt engineering has its limitations in reflecting unique styles, especially those not present in the pre-training data.
@@
To achieve more effective style-driven generation, previous research has explored methods such as fine-tuning [47, 53, 32], image variation [62] and editing [15]. These approaches commonly use style reference images and their accompanying text prompts to guide the generation process. However, they blindly apply the same objective functions and noise level distributions used for pre-training—originally optimized for object-centric benchmarks [12, 22]—without acknowledging that style images emphasize different visual factors such as color schemes, layout, and brushstrokes, rather than simply depicting specific objects. Consequently, even with numerous style-driven generation studies [53, 32, 62, 46, 15], we observe prevalent failure cases when fine-tuning diffusion models on style references. Addressing these failures sometimes requires two-stage training with heavy human feedback [53].
In this paper, we address these limitations by introducing the Style-friendly SNR sampler, a method that ensures capturing novel styles during fine-tuning. Our approach is motivated by two key observations: 1) diffusion models [9, 31] struggle to learn new styles, and 2) styles emerge at higher noise levels. These observations reveal that standard fine-tuning procedures often devote a large portion of computation toward lower noise levels, which are less relevant for learning novel styles. Building upon these observations, we propose reallocating fine-tuning computation on higher noise levels, where essential style attributes emerge.
@@
Our method, focusing on the reallocation of noise level sampling during fine-tuning, enhances the fidelity of style-driven generation, capturing the novel styles indicated by text prompts while preserving the capability of depicting the desired content. Moreover, we unveil the key components that explain why these models excel at learning object-centric concepts but struggle with styles, providing deeper insights into the diffusion process for style-driven generation. Ultimately, our approach facilitates the creation of style templates from reference images, which can be easily shared and utilized to practitioners for content creation, expanding the capabilities of diffusion models.
@@ 2 Training Diffusion Models
Figure 2: Probability distribution of Log-SNR. We bias the distribution towards the shaded region where style features emerge.
2.1 Diffusion Process and SNR Formulation
@@
where $\alpha_{t}$ and $\sigma_{t}$ are predefined noise schedules, and $\epsilon\sim\mathcal{N}(0,I)$ represents standard Gaussian noise.
Recent state-of-the-art flow matching frameworks, such as Stable Diffusion 3 (SD3) [9] and FLUX [31], utilize the noise schedule from rectified flow [35, 36], where $\alpha_{t}=1-t$ and $\sigma_{t}=t$, with $t$ varying continuously in the range $[0,1]$. This choice is effective due to straight diffusion trajectories.
Instead of parameterizing the diffusion process using the timestep $t$, Kingma et al. [27, 26] characterize the noise level using the log signal-to-noise ratio (log-SNR), which offers a more intuitive measure of the noise at each step:
@@
In flow matching framework, the log-SNR is simplified as $\lambda_{t}=2\log\left(\frac{1-t}{t}\right)$ using timestep $t$.
2.2 Diffusion Training and SNR Sampling
Recent flow matching frameworks predict the velocity field $v_{\theta}(x_{t},t)$ by minimizing the following training objective:
$\mathcal{L}_{\text{FM}}(x_{0})\;=\;\mathbb{E}_{t\sim p(t)}\Big{[}\big{\|}\epsilon-x_{0}-v_{\theta}(x_{t},t)\big{\|}^{2}\Big{]},$ (3)
where $p(t)$ is the timestep sampling distribution, $\epsilon-x_{0}$ is the target velocity derived from the forward process in Eq. 1. The representative text-to-image models utilizing flow matching formulation, such as SD3 [9] and FLUX [31], introduce a SNR sampler for training. This sampler samples the logit of $t$, defined as $\log\left(\frac{t}{1-t}\right)$ from $\mathcal{N}(\mu,\sigma^{2})$, where the parameters $\mu$ and $\sigma$ are chosen as 00 and $1$ to optimize CLIP [42] and FID scores [16] on COCO-2014 validation set [34].
In addition, they propose shifting timestep $t$ to $t_{\text{new}}$ by $k$ for high resolution training:
@@ 2.2 Diffusion Training and SNR Sampling
$\lambda_{t_{\text{new}}}=2\log\left(\frac{1-t_{\text{new}}}{t_{\text{new}}}\right)=\lambda_{t}-2\log k,$ (5)
where $k$ is defined as $3$. Following the above formulation, resulting log-SNR sampling distribution $p(\lambda_{t})$ in training time is represented as $\mathcal{N}(-2\log 3,2^{2})$, as visualized by the yellow curve in Fig. 2. This curve demonstrates that pre-training SD3 and FLUX focus on particular noise levels.
@@ 3.1 Observations
Figure 3: Fine-tuning capability. (a) While FLUX succeeds in learning objects, (b) it struggles to capture styles, demonstrating that learning novel objects and styles requires distinct strategies. (c) We enable FLUX to learn styles. References are shown in the red insert box.
Figure 4: Prompt switching during generation. $\lambda_{t}$ indicates log-SNR. The bar graphs above each image represent the denoising steps, illustrating when each prompt is applied and at what point the prompt switch occurs. The style prompts are ‘minimalist flat round logo’, ‘sticker’, ‘detailed pen and ink drawing’, and ‘cartoon’. Styles emerge in the initial 10% of denoising steps; therefore, (c) and (f) fail to capture target styles. In contrast, omitting style prompts in later steps (d,e) still preserves styles well, similar to the fully styled baseline (a). (g) and (h) quantify these observations, showing the average CLIP similarity across 5 prompts and 5 styles when omitting (g) or including (h) the style prompt in earlier steps. Here, we use FLUX with 28 inference steps.
@@
We begin by examining the fine-tuning capabilities of current text-to-image diffusion models, which have primarily been studied for object-driven generation [47] from reference images. Fine-tuning is typically guided by accompanying text prompts that specify the desired objects or styles. As shown in Fig. 3a, these methods often excel at producing high-quality object-driven images. However, we observe poorer results when they are applied to replicate distinctive color schemes, illumination, or brushwork—particularly when using the same SNR sampler from pre-training [9]. In Fig. 3b, when provided with a “glowing” style reference, the FLUX model only applies glowing effects to specific object details (such as the fur of a sloth), neglecting broader stylistic elements like the dark background and blue lighting of the original reference. Similarly, with a Van Gogh oil painting style reference, the model manages to replicate the blue color tone but fails to accurately reproduce distinctive brushstroke characteristics.
@@
To better understand why diffusion models struggle to capture new styles, we investigate at which noise levels stylistic features emerge during generation using a pre-trained FLUX model [31]. Specifically, we switch from a prompt without style descriptions ($y_{\text{w/o style}}$) to one including style descriptions ($y_{\text{w/ style}}$) at different points in the denoising process. Omitting the style prompt in the initial 10% of generation steps significantly reduces style alignment (Fig. 4c), as quantitatively confirmed by the low CLIP similarity scores when initial denoising steps omit style information (Fig. 4g). Conversely, omitting the style prompt at later denoising steps minimally affects style alignment (Fig. 4d,e), which is also supported quantitatively in Fig. 4h. These results demonstrate that styles are predominantly determined at early denoising steps, corresponding to intervals with higher noise levels (low log-SNR $\lambda_{t}$ values).
@@ 3.2 Style-Friendly SNR Sampler
Our earlier observations show that styles primarily emerge during early denoising steps, characterized by higher noise levels (lower log-SNR values). However, existing fine-tuning methods use the SNR sampler from pre-training, optimized mainly for object-centric benchmarks [22, 12], as indicated by the yellow curve in Fig. 2. Consequently, standard fine-tuning procedures place insufficient emphasis on noise levels crucial for capturing styles, failing to achieve alignment with reference styles.
$\lambda_{t}\sim\mathcal{N}(\mu_{\text{low}},\sigma_{\text{large}}^{2}),$ (6)
[…]
While timestep shifting in [9] weakly biases the noise level distribution as in Eq. 5, our Style-friendly SNR sampler aggressively targets the high-noise regions critical for capturing style. Setting the mean to $\mu_{\text{low}}=-6$ and $\sigma_{\text{large}}\geq 2$ targets the shaded region of Fig. 2, substantially improving style alignment across a variety of reference styles.
3.3 Trainable Parameters of MM-DiT
We fine-tune both FLUX-dev [31] and SD3.5 [9, 59] by training LoRA [20] adapters on specific layers to capture new styles. Multi-Modal Diffusion Transformer (MM-DiT) [9], the core architecture for both FLUX-dev and SD3.5 comprises dual-stream transformer blocks with separate parameters for text tokens and image tokens, which interact through joint attention mechanisms. To effectively learn the stylistic features encompassing both visual and linguistic characteristics, we train LoRA on the attention layers of both modalities. Additionally, FLUX includes single-stream blocks [6] that handle both modalities simultaneously with attention mechanisms and projection layers that skip this attention, to which we also apply LoRA. This targeted fine-tuning achieves high style-alignment without training the entire network, providing a parameter-efficient method for fine-tuning MM-DiT.
Figure 5: Effect of varying $\bm{\mu}$ and $\bm{\sigma}$. Diffusion models start to capture the reference glowing style when $\mu$ is lower and $\sigma$ is larger. The prompt is ‘a hybrid creature that is a mix of a waffle and a hippopotamus, in glowing style’. Samples are generated with the same seed.
@@ 3.3 Trainable Parameters of MM-DiT
Figure 6: SNR sampler analysis. DINO similarities of varying SNR sampler parameters with FLUX and SD3.5-8B. Dotted lines in (c) indicate results of SD3 sampler [9]. Unless specified, we use $\mu=-6$, $\sigma=2$, and rank 32. CLIP scores are shown in the Appendix.
@@ 4 Experiments
Figure 7: Qualitative comparison. We show ‘A fluffy baby sloth with a knitted hat trying to figure out a laptop’, ‘A christmas tree’, and ‘An espresso machine’ in various styles. Target style prompts are shown above. Our Style-friendly SNR sampler effectively captures the styles indicated by both reference images and their accompanying text prompts. Fine-tuning methods (SD3 sampler and DCO) often miss stylistic nuances due to insufficient focus on the relevant noise levels or overly strong regularization.
We compare our method against both fine-tuning and non-fine-tuning approaches for style-driven generation.
  • Fine-tuning: SD3 sampler [9], Direct Consistency Optimization (DCO) [32], and StyleDrop [53] are selected as representative fine-tuning methods. The SD3 sampler utilizes a flow matching loss [35, 36] with timestep shifting in Eq. 5, while DCO is based on reinforcement learning [43] and employs a regularized loss. For StyleDrop, due to the lack of official open-source implementations, we adopt the unofficial implementation provided by [1].
  • Image variation: IP-Adapter [62] and RB-Modulation [46] generate image variations using CLIP [42] or CSD [54] image embeddings respectively. IP-Adapter reconstructs images to best reflect the information contained in the CLIP image embedding, whereas RB-Modulation applies gradient guidance at test time to maximize the similarity score evaluated by CSD.
  • Editing: Style-Aligned [15] is an editing-based method that shares and manually uplifts self-attention weights to enforce the reference style consistency.
  • Detailed prompt: We name this baseline GPT-4o Prompt, where we utilize GPT-4o [23] to generate detailed style descriptions from reference images and apply these in text-to-image generation without fine-tuning.
Following StyleDrop [53], we select 18 reference styles as our fine-tuning targets. For each style, we use 23 evaluation prompts, generating 2 images per prompt, which results in a total of 828 images per experiment. We fine-tune FLUX-dev [31] and SD3.5-8B [9, 59] with LoRA [20] at rank 32 (except where noted for rank ablation). Following [53, 15] we measure style alignment via human evaluation, DINO ViT-S/16 [3], and CLIP ViT-B/32 [42] image similarity (CLIP-I), and alignment to the text prompts via CLIP text-image similarity (CLIP-T). Additional implementation details are in the Appendix.